OpenStudy (anonymous):

SOLVE FOR X: 3/4x = 9

OpenStudy (anonymous):

3/4x=9 3x=36 x=12

OpenStudy (anonymous):

3=9x4x (36x) 3=36x 35/3=x x=12

OpenStudy (anonymous):

3x=9*4=36 3x=36/3=12